OpenStudy (anonymous):

What is the simplified form of the quantity of x plus 8, all over the quantity of x plus 2 + the quantity of x minus 3, all over 4? the quantity of x squared plus 5x minus 24, all over 4 times the quantity of x plus 2 the quantity of x squared plus 3x plus 26, all over 4 times the quantity of x plus 2 the quantity of x squared plus 3x plus 26, all over x plus 6 the quantity of x squared plus 5x minus 24, all over 4 times the quantity of x plus 6 I think the answer is either A or D. Can someone help me determine which one?

OpenStudy (anonymous):



OpenStudy (anonymous):

@bibby i was gonna tag you but i figured youd see this anyways lol

OpenStudy (bibby):

lol, so we want to get them into a common denominator again \(\dfrac{x+8}{x+2}+\dfrac{x-3}{4}=\dfrac{(x+8)(4)}{(x+2)(4)}+\dfrac{(x-3)(x+2)}{(4)(x+2)}\)

OpenStudy (bibby):

\(\dfrac{4x+32+x^2-x-6}{4x+8}=\dfrac{x^2+3x+26}{4x+8}\)

OpenStudy (bibby):

think I messed up somewhere

OpenStudy (anonymous):

uh oh. Its okay.

OpenStudy (bibby):

f